# Differentiation Through Black-Box Quadratic Programming Solvers

### v3

@@ Appendix B Proof of Theorem 1
$\begin{bmatrix}\partial_{\theta}z^{*}\\ \partial_{\theta}\lambda^{*}\\ \partial_{\theta}\mu_{J}^{*}\end{bmatrix}=-\begin{bmatrix}P&A^{T}&C_{J}^{T}\\ A&0&0\\ C_{J}&0&0\\ \end{bmatrix}^{-1}\Biggl{(}\begin{bmatrix}\partial_{\theta}P&\partial_{\theta}A^{T}&\partial_{\theta}C_{J}^{T}\\ \partial_{\theta}A&0&0\\ \partial_{\theta}C_{J}&0&0\\ \end{bmatrix}\begin{bmatrix}z^{*}\\ \lambda^{*}\\ \mu_{J}^{*}\end{bmatrix}-\begin{bmatrix}-\partial_{\theta}q\\ \partial_{\theta}b\\ \partial_{\theta}d_{J}\end{bmatrix}\Biggr{)}$. (12)
@@
which are exactly the KKT conditions of the equality-constrained reduced QP (5). Uniqueness of solution then implies that (1) and (5) are pointwise equivalent at $\theta$. Moreover, since $P,q,A,b,C,d$ are smoothly parameterized by $\theta$, the Basic Sensitivity Theorem [Fiacco, 1976] asserts that the primal-dual solution $\zeta^{*}(\theta)$ for (1) is a differentiable function of $\theta$ in a neighborhood of $\theta$, defined implicitly through the KKT’s equality conditions. Furthermore, the active set $J(\theta)$ is fixed in this neighborhood, therefore (1) and (5) are locally equivalent.
@@
yielding (6). Moreover, since $J(\theta)$ is locally constant, the Basic Sensitivity Theorem establishes that $\zeta_{J}^{*}(\theta)$ is differentiable. Using the formal derivative of the matrix inverse [Magnus & Neudecker, 1988; Petersen & Pedersen, 2008] we explicitly differentiate (16) to obtain
@@ Appendix C Backpropagation
Like other differentiable QP layers implemented within automatic differentiation frameworks such as PyTorch [Paszke et al., 2019], we do not directly return the derivative $\partial_{\theta}\zeta^{*}$. This is in part because dQP directly receives the QP parameters $P,q,A,b,C,d$ and not $\theta$, and so in backpropogation we are not concerned with $\theta$. This is rather accounted for in the next step outside dQP, usually by automatic differentiation. Secondly, backpropagation requires that we compute a so-called Jacobian-vector product which are products of the Jacobians with an “incoming” gradient of a quantity or loss $\ell$ that depends on $\zeta^{*}$. This requires less computation and does not require the formation of a 3-tensor. Since $\zeta_{J}^{*}=K_{J}^{-1}v_{J}$ is a formal matrix-vector multiplication, the Jacobian-vector product is well-known,
@@
similar to OptNet [Amos & Kolter, 2017a]. We note that the gradient with respect to $P$ is constrained to lie within the subspace of symmetric matrices. Similarly, if the matrices $P,A,C$ are sparse, then we project the gradient to lie within the non-zero entries, which can be implemented efficiently in Equations 20. Although the above argument is for a scalar loss $\ell$, the same approach is naturally adapted if $\zeta^{*}$ is mapped to a vector in the immediate next layer.
@@ Appendix D Implementation Details
Choosing a solver. Since our work enables users to choose any QP solver as the front-end for their differentiable QP applications, we include a simple diagnostic tool for quantitatively measuring solver performance. We present an example result in Figure 6 for the cross geometry experiment in section 4, finding PIQP, OSQP, and QPALM to be the most efficient. For this reason, we choose PIQP in the geometry experiments. We also include tools for checking the solution and gradient accuracy.
Figure 6: Evaluating the best QP solver for the cross geometry problem using our diagnostic tool. The solution tolerance regimes are varied between $\epsilon_{\textrm{abs}}=10^{-8},10^{-5},10^{-2}$.
@@
Convexity and Feasibility. Two key assumptions of our method are strict convexity and feasibility. However, these are often violated in practice. We include optional checks that $P$ is symmetric positive definite. On the other hand, we do not perform any special handling for infeasibility – a limitation of our method compared to, for example, QPLayer [Bambade et al., 2024].
@@
QP Solvers. Throughout this work, we use a number of QP solvers available in qpsolvers including Clarabel [Goulart & Chen, 2024], DAQP [Arnström et al., 2022], Gurobi [Gurobi Optimization, LLC, 2024], HiGHS [Huangfu & Hall, 2018], HPIPM [Frison & Diehl, 2020], MOSEK [Andersen & Andersen, 2000], OSQP [Stellato et al., 2020], PIQP [Schwan et al., 2023], ProxQP [Bambade et al., 2023], QPALM [Hermans et al., 2019], qpSWIFT [Pandala et al., 2019], quadprog [Goldfarb & Idnani, 1983], and SCS [O’Donoghue et al., 2016].
@@ E.1 Performance Evaluation
Since SCQPTH does not support equality constraints, we convert them into an corresponding set of inequality constraints.
@@ E.1.1 Projection onto the probability simplex
For projection onto the probability simplex, as formulated in $P_{1}$, equation (8), we set $x\in\mathbb{R}^{n}$ with $x_{i}\sim\mathcal{N}(0,1)$. The dataset, with 500 problems, has dimensions $n\in$ {10, 20, 50, 100, 220, 450, 1000, 2100, 4600, 10000, 100000}. For $n\leq 4600$, each dimension contains 50 problems and 25 problems for $n>4600$. Gurobi is chosen as dQP’s forward solver. Figure 1 shows the median performance within the $1/4$ and $3/4$ quantiles for each dimension. SCQPTH failed for all problems with $n>50$. The statistics in Table 5 show that we outperform competing methods for differentiable QP in both forward and backward times.
@@ E.2 Bi-Level Geometry Optimization
We include a supplementary example in Figure 9 which illustrates the boundary conditions (inequality constraints in (10)) of [Kovalsky et al., 2015] using blue arrows (corresponding to applying the Laplacian to the output vertices) and cones at points where the shape is locally non-convex. If the arrows fall outside the cones as shown in (b), then the map is non-invertible; if instead they lie strictly inside the cones then the map is invertible as shown in (c). The process is also flexible: by adding a regularizing term $\lambda\|\frac{M}{\|M\|_{F}}-\frac{M_{c}}{\|M_{c}\|_{F}}\|_{\infty}$ that measures the distance to the combinatorial Laplacian $M_{c}$ shown in (d), we can enhance map quality in the sense that the triangles change their shape less with respect to the input mesh (a).
@@
The upper-level loss for the bi-level cross experiment is shown in Figure 10(a) where the unregularized loss is driven to the desired tolerance, accompanied by sudden changes in the active set as the dual variables are driven to zero. We terminate the optimization at convergence, once all constraints are inactive to guarantee an injective map. For the regularized optimization (Figure 10(b)), we choose the regularization hyper-parameter to be $\lambda=10$ after sample testing. Despite the regularization, the dual loss can still be driven to 0 to reach an injective map.
@@
Throughout the geometry experiments, we use the same solution tolerance $\epsilon_{\textrm{abs}}=10^{-5}$ and active tolerance $\epsilon_{J}=10^{-4}$ with the forward solver PIQP as determined by our tool for choosing the best solver described in Appendix D. For the upper-level optimization, we use the Adam optimizer with learning rate $10^{-2}$ [Kingma & Ba, 2017]. We initialize the bi-level optimization with $M_{c}$.
@@
The cross (Figure 9) and ant (Figure 5(a)) meshes and boundary constraints are obtained from the datasets in [Du et al., 2020]. We create the synthetic mesh refinement example in Figure 5(b) by perturbing the corner of a square mesh.

### v4

@@ Appendix B Proof of Theorem 1
[…]
@@
which are exactly the KKT conditions of the equality-constrained reduced QP (5). Uniqueness of solution then implies that (1) and (5) are pointwise equivalent at $\theta$. Beyond this, since $P,q,A,b,C,d$ are smoothly parameterized by $\theta$, the Basic Sensitivity Theorem [46] asserts that the primal-dual solution $\zeta^{*}(\theta)$ for (1) is a differentiable function of $\theta$ in a neighborhood of $\theta$, defined implicitly through the KKT’s equality conditions, and that the active set $J(\theta)$ is fixed in this neighborhood. Thus, (1) and (5) are locally equivalent in a neighborhood of the parameter $\theta$.
@@
yielding (6). Moreover, since $J(\theta)$ is constant in a local neighborhood, the Basic Sensitivity Theorem establishes that $\zeta_{J}^{*}(\theta)$ is differentiable. Using the derivative of the matrix inverse [78, 91], we explicitly differentiate (16) to obtain
@@ Appendix C Backpropagation
Like other differentiable QP layers implemented within automatic differentiation frameworks such as PyTorch [88], we do not directly return the derivative $\partial_{\theta}\zeta^{*}$. This is in part because dQP directly receives the QP parameters $P,q,A,b,C,d$ and not $\theta$, and so in backpropogation we are not concerned with $\theta$. This is rather accounted for in the next step outside dQP, usually by automatic differentiation. Secondly, backpropagation requires that we compute a so-called Jacobian-vector product which are products of the Jacobians with an “incoming” gradient of a quantity or loss $\ell$ that depends on $\zeta^{*}$. This requires less computation and does not require the formation of a 3-tensor. Since $\zeta_{J}^{*}=K_{J}^{-1}v_{J}$ is a formal matrix-vector multiplication, the Jacobian-vector product is well-known,
@@
similar to OptNet [5]. We note that the gradient with respect to $P$ is constrained to lie within the subspace of symmetric matrices. Similarly, if the matrices $P,A,C$ are sparse, then we project the gradient to lie within the non-zero entries, which can be implemented efficiently in Equations 20. Although the above argument is for a scalar loss $\ell$, the same approach is naturally adapted if $\zeta^{*}$ is mapped to a vector in the immediate next layer.
@@ Appendix D Implementation Details
Choosing a solver. Since our work enables users to choose any QP solver as the front-end for their differentiable QP applications, we include a simple diagnostic tool for quantitatively measuring solver performance. We present an example result in Figure 6 for the cross geometry experiment in Figure 9, finding PIQP, OSQP, and QPALM to be the most efficient. For this reason, we choose PIQP in the geometry experiments. We also include tools for checking the solution and gradient accuracy.
Figure 6: Evaluating the best QP solver for the cross geometry problem (Figure 9) using our diagnostic tool. The solution tolerance regimes are varied between $\epsilon_{\textrm{abs}}=10^{-8},10^{-5},10^{-2}$.
@@
Convexity and Feasibility. Two key assumptions of our method are strict convexity and feasibility. However, these are often violated in practice. We include optional checks that $P$ is symmetric positive definite. On the other hand, we do not perform any special handling for infeasibility – a limitation of our method compared to, for example, QPLayer [17].
@@
QP Solvers. Throughout this work, we use a number of QP solvers available in qpsolvers including Clarabel [55], DAQP [13], Gurobi [58], HiGHS [64], HPIPM [51], MOSEK [10], OSQP [108], PIQP [104], ProxQP [16], QPALM [60], qpSWIFT [87], quadprog [54], and SCS [85].
@@ E.1 Performance Evaluation
Since SCQPTH does not support equality constraints, we convert them into a corresponding set of inequality constraints.
@@ E.1.1 Projection onto the probability simplex
For projection onto the probability simplex, as formulated in $P_{1}$, equation (8), we set $x\in\mathbb{R}^{n}$ with $x_{i}\sim\mathcal{N}(0,1)$ drawn randomly from a standard normal distribution. The dataset, with 500 problems, has dimensions $n\in$ {10, 20, 50, 100, 220, 450, 1000, 2100, 4600, 10000, 100000}. For $n\leq 4600$, each dimension contains 50 problems and 25 problems for $n>4600$. Gurobi is chosen as dQP’s forward solver. Figure 1 shows the median performance within the $1/4$ and $3/4$ quantiles for each dimension. SCQPTH failed for all problems with $n>50$. The statistics in Table 5 show that dQP outperforms competing methods for differentiable QP in both forward and backward times.
@@ E.2 Bi-Level Geometry Optimization
We include a supplementary example in Figure 9 which illustrates the boundary conditions (inequality constraints in (10)) of [68] using blue arrows (corresponding to applying the Laplacian to the output vertices) and cones at points where the shape is locally non-convex. If the arrows fall outside the cones as shown in (b), then the map is non-invertible; if instead they lie strictly inside the cones then the map is invertible as shown in (c). The process is also flexible: by adding a regularizing term $\lambda\|\frac{M}{\|M\|_{F}}-\frac{M_{c}}{\|M_{c}\|_{F}}\|_{\infty}$ that measures the distance to the combinatorial Laplacian $M_{c}$ shown in (d), we can enhance map quality in the sense that the triangles change their shape less with respect to the input mesh (a).
@@
The upper-level loss for the bi-level cross experiment is shown in Figure 10(a) where the unregularized loss is driven to the desired tolerance, accompanied by sudden changes in the active set as the dual variables are driven to zero. We terminate the optimization at convergence, once all constraints are inactive to guarantee an injective map. For the regularized optimization (Figure 10(b)), we choose the regularization hyperparameter to be $\lambda=10$ after sample testing. Despite the regularization, the dual loss can still be driven to 0 to reach an injective map.
@@
Throughout the geometry experiments, we use the same solution tolerance $\epsilon_{\textrm{abs}}=10^{-5}$ and active tolerance $\epsilon_{J}=10^{-4}$ with the forward solver PIQP as determined by our tool for choosing the best solver described in Appendix D. For the upper-level optimization, we use the Adam optimizer with learning rate $10^{-2}$ [67]. We initialize the bi-level optimization with $M_{c}$.
@@
The cross (Figure 9) and ant (Figure 5(a)) meshes and boundary constraints are obtained from the datasets in [41]. We create the synthetic mesh refinement example in Figure 5(b) by perturbing the corner of a square mesh.